# Enhancing and Assessing Instruction-Following with Fine-Grained Instruction Variants

### v1

@@ Abstract
The effective alignment of Large Language Models (LLMs) with precise instructions is essential for their application in diverse real-world scenarios. Current methods focus on enhancing the diversity and complexity of training and evaluation samples, yet they fall short in accurately assessing LLMs’ ability to follow similar instruction variants. We introduce an effective data augmentation technique that decomposes complex instructions into simpler sub-components, modifies these, and reconstructs them into new variants, thereby preserves the original instruction’s context and complexity while introducing variability, which is critical for training and evaluating LLMs’ instruction-following precision. We developed the DeMoRecon dataset using this method to both fine-tune and evaluate LLMs. Our findings show that LLMs fine-tuned with DeMoRecon will gain significant performance boost on both ours and commonly used instructions-following benchmarks.
@@ 1 Introduction
To address these challenges, we propose an effective method that improves LLMs’ instruction-following capabilities by Decompose each instruction into sub-instructions, Modifying them, and then Reconstruct these into new variants. This process preserves the original context and difficulty while increasing variability. We have developed the DeMoRecon datasets to both enhance and evaluate these capabilities. The DeMoRecon-Aug and DeMoRecon-Aug-DPO datasets uses direct responses from GPT-4 based on the augmented instructions for SFT and DPO, whereas DeMoRecon-Ref and DeMoRecon-Ref-DPO incorporates responses revised from the original dataset to maintain closeness to the initial instructions. Additionally, the DeMoRecon-Eval dataset, which includes 170 instructions and their augmentations, leverages GPT-4 to assess the accuracy of following very similar instruction variants.
Our assessment of the DeMoRecon dataset’s effectiveness involves applying instruction tuning and alignment tuning to fine-tune LLMs and evaluating their performance across DeMoRecon-Eval and established benchmarks like IFEval, InfoBench, and FollowBench. The results demonstrate notable enhancements in the precision with which LLMs follow instructions. Key contributions of our work include:
  • We introduce a data augmentation method that breaks down complex instructions into simpler sub-components, modifies them, and reconstructs them into new variants. This preserves the original instruction’s difficulty and context while increasing variability, thereby improving LLMs’ ability to detect nuanced differences in commands.
  • We develop the DeMoRecon dataset, which includes five components (DeMoRecon-Aug, DeMoRecon-Ref, DeMoRecon-Aug-DPO, DeMoRecon-Ref-DPO, and DeMoRecon-Eval), targeting both enhanced training and rigorous evaluation of LLMs’ instruction-following abilities. This dataset is specifically designed to test LLMs on subtle variations within instructions, providing a refined tool for assessing nuanced comprehension and execution.
  • Through comprehensive testing, we show that LLMs trained with DeMoRecon outperform those trained with traditional benchmarks in nuanced instruction-following tasks. Our findings, validated using several leading LLMs including GPT-4, establish DeMoRecon as a critical benchmark for future assessments in this field. We have made the training data and source code fully open-source to support ongoing research.
@@ 2 Approach
Figure 1: An illustration of the proposed method of constructing DeMoRecon.
In this section, we describe the process of constructing DeMoRecon and detail its advantages. Figure 1 provides an illustration of our proposed instruction augmentation method, while Table 1 presents the dataset statistics for DeMoRecon. We utilize gpt-4-0125-previewhttps://platform.openai.com/docs/models for all tasks involving large language models (LLMs). The prompt templates designed and used in our experiments are included in the Appendix.
@@ 2.1 Seed Preparation
To adequately encompass instructions of varying difficulty levels, we elected to source initial seeds from instructions developed by Xu et al. (2023) using Evol-Instructhttps://huggingface.co/datasets/WizardLMTeam/WizardLM_evol_instruct_V2_196k. We began by randomly sampling 2,000 instructions from the official dataset. Subsequently, GPT-4 was employed to exclude any instructions lacking sufficient background information. For instance, some instructions requested a summary of a passage that was not provided. After this filtering process, we established a final valid seed set, $\mathbf{S}=\{P_{i}\}_{1\leq i\leq N}$, where $N=1,773$ represents the number of valid instructions encompassing both easy and challenging topics. These seeds were then utilized to construct our DeMoRecon.
@@ 2.2 Instruction Augmentation
Reconstruction. After modifying the sub-instruction, we reconstruct $P_{i}$ into $\bar{\mathbf{P}}_{i}^{A}=\{P_{i}\}\cup\{\bar{P}_{i}^{j}\}_{1\leq j\leq M_{i}}$, where $\bar{P}_{i}^{j}$ is formed by prompting GPT-4 with the original facts, $\mathbf{F}_{i}$, and the newly modified sub-instruction set, $\bar{\mathbf{I}}_{i}^{j}$. Specifically, in this step, we provide GPT-4 with the original instruction $P_{i}$, its sub-instructions $\mathbf{I}_{i}$, its background facts $\mathbf{F}_{i}$, and the sub-instructions $\bar{\mathbf{I}}_{i}^{j}$ where one of them are modified. We instruct GPT-4 it to revise only the necessary parts of the original instruction $P_{i}$ to accommodate the changes in sub-instructions. This approach is designed to keep most of the original prompt unchanged, making the modified sub-instruction the sole variant between the original and the new instructions.
@@ 2.3 Response Collection
For evaluation purposes, we select a subset of 170 instructions from $S$ that contain four sub-instructions, and use their augmented variants to form the test set. This limitation is imposed because GPT-4 is more prone to errors when processing overly complex instructions.
For DPO training, we construct preference pairs for preparation. For each instruction pair $(P_{i},P_{j})_{P_{i},P_{j}\in\bar{\mathbf{P}}_{i}^{A}}$, we label the response of $P_{i}$ as “good” and the response of $P_{j}$ as “bad” to create a DPO training sample. To reduce the size of the training data, we keep at most 5 pairs for each seed instructions by random sampling. We perform such process on both DeMoRecon-Aug and DeMoRecon-Ref to construct DeMoRecon-Aug-DPO and DeMoRecon-Ref-DPO.
Consequently, we construct five versions of DeMoRecon: 1) DeMoRecon-Eval, comprising 680 instructions designated for assessing the instruction-following capabilities of LLMs on fine-grained instruction variants; 2) DeMoRecon-Aug, which includes the remaining original instructions along with their variants and the responses collected via direct prompting of GPT-4; 3) DeMoRecon-Ref, which utilizes the original instructions and their responses as references to gather responses for the instruction variants in DeMoRecon-Aug; 4) DeMoRecon-Aug-DPO, the DPO data for DeMoRecon-Aug and 5) DeMoRecon-Ref-DPO, the DPO data for DeMoRecon-Ref.
@@ 2.4 Statistic and Analysis
Table 1: The Data statistic of DeMoRecon.
DeMoRecon- Eval Aug Ref Aug-DPO Ref-DPO Number of Instructions 680 8,705 8,705 6,445 6,445 Instruction Average Token 74 126 126 125 125 Response Average Token 466 656 631 - - $U$ of Response - 0.036 0.064 - -
Figure 2: The tSNE plots of the semantic embeddings of DeMoRecon. The upper plot is for the original and their augmented instructions. The lower plot represents the responses of the original instruction, augmented instruction and augmented instruction using reference.
Figure 2 displays the t-SNE 2D projection of the semantic embeddings of the instructions and responses in DeMoRecon-Aug and DeMoRecon-Ref. In the top plot, gray points represent the original instructions in $\mathbf{S}$, and orange points represent their augmented instructions. This proximity on the plot indicates that the semantic similarity between the original and augmented instructions is high, which benefits LLMs during fine-tuning by exposing them to instructions with minor variations.
In the lower plot, the blue points represent the responses to the augmented instructions using references, while the orange points represent the responses without references. The blue points largely overlap with the gray points, which are the responses to the original instructions, indicating that the responses using references are more similar to the original responses and thus align with our design objectives. The orange points, in contrast, are more dispersed, suggesting that responses without references diverge more from the original responses. This distribution demonstrates that with minimal semantic changes, LLMs can learn to generate consistent responses to similar instructions.
Following the methodology in Guo et al. (2023), we calculate the variance of the nearest neighbor distances for the responses of DeMoRecon-Aug and DeMoRecon-Ref to quantify their dispersion:
@@
Here, $U$ represents the distribution uniformity, $Q$ denotes the number of responses, and $\mu$ indicates the average Euclidean distance between the semantic embedding $e_{k}$ of the $k$-th response and the context embedding $e^{\text{NN}}_{k}$ of its nearest neighbor. A larger value of $U$ implies that the responses are less uniform and thus more similar to each other. Given our goal to make the responses as similar to the original responses as possible, we expect to observe higher values of $U$ on the DeMoRecon-Ref dataset.
The calculated results, presented in Table 1, show that the distribution uniformity of the responses in DeMoRecon-Ref is 0.064, while in DeMoRecon-Aug, it is 0.035. This difference further supports the observation that the responses obtained using references are more similar to the original responses than those obtained directly from DeMoRecon-Aug.
@@ 3 Experiments
In this section, we detail the experiments conducted with LLMs fine-tuned using DeMoRecon. We assess the accuracy of instruction-following based on evaluations from GPT-4 and rule-based (IFEval and partial FollowBench) assessments under greedy generation settings. The performance is reported across DeMoRecon-Eval and three other well-known instruction-following benchmarks.
@@ 3.1 Baselines
These models, available in their officially released instruction-tuned versions, are employed to precisely assess their instruction-following capabilities across various instruction variants and to explore the potential for further enhancements in this area. The specific releases utilized are detailed in Table 2 and can be accessed on Hugging Face§§§https://huggingface.co.
Note that our analysis focuses on comparing the performance between the baseline instruction-tuned models and those versions additionally fine-tuned with DeMoRecon set. This approach ensures a fair comparison by directly assessing and potentially enhancing the instruction-following capabilities of already instruction-tuned LLMs. Unlike other research (Sun et al., 2024) that primarily aims to improve instruction-following from pre-trained LLMs—a process that can be both costly and time-intensive—our study also addresses recent updates in model versions. The studies often rely on gpt-4-0314, a version that has recently been deprecatedhttps://platform.openai.com/docs/deprecations. For all GPT-4 based evaluations, we use gpt-4-0125-preview.
@@ 3.2 Benchmarks
Besides DeMoRecon-Eval, we extend our performance comparison to include three of the most popular instruction-following benchmarks, further demonstrating the efficacy of our methods.
@@ 3.3 Experimental Settings
We employed both Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) on base chat models. Specifically, for SFT, we directly fine-tuned base chat LLMs on DeMoRecon-Aug and DeMoRecon-Ref. For comparison, we also fine-tuned these models using only the seed instructions and their respective responses, with these results labeled as “plain” in Table 2.
For DPO, we integrated SFT and DPO losses to fine-tune the LLMs, addressing both preference alignment and response generation, which are crucial for effective instruction-following. DPO focuses on enhancing the model’s preference for better adherence to instructions, while SFT ensures the generation of correct, precise, and high-quality responses. This combined approach has proven effective in our ablation studies. Specifically, for each preference sample, we use the “good” response to supervise the fine-tuning of LLMs alongside DPO-based learning.
@@ 3.4 Experimental Results
Table 2: Experimental results on DeMoRecon-Eval and other instruction-following benchmarks. “Aug.” represent DeMoRecon-Aug and DeMoRecon-Aug-DPO, “Ref.” represents DeMoRecon-Ref and DeMoRecon-Ref-DPO. The performance of the original chat LLMs are recorded under “Ori.”.
Model DeMoRecon-Eval Ori. SFT DPO+SFT Plain Aug. Ref. Aug. Ref. GPT-4 95.00 - - - - - GPT-4-Ref 91.03 - - - - - Llama2-7B-Chat 54.71 56.32 59.12 59.26 60.29 60.44 Qwen1.5-7B-Chat 61.91 71.03 70.74 69.56 72.06 70.59 Mistral-7B-Instruct 64.41 66.32 67.79 68.24 67.79 71.18 Llama2-13B-Chat 58.82 62.56 64.12 63.38 64.41 65.15 Qwen1.5-14B-Chat 73.53 72.06 76.32 73.97 79.26 81.62 Orion-14B-Chat 56.62 62.21 67.06 66.03 68.82 68.38 FollowBench (average HSR) Llama2-7B-Chat 43.10 44.49 44.02 43.28 45.56 46.41 Qwen1.5-7B-Chat 53.00 53.36 51.09 52.77 51.36 54.86 Mistral-7B-Instruct 42.55 46.56 47.18 46.37 49.63 47.68 Llama2-13B-Chat 44.94 46.56 45.85 46.79 47.23 48.07 Qwen1.5-14B-Chat 55.32 50.09 52.05 51.93 58.02 56.25 Orion-14B-Chat 44.12 48.81 49.17 48.71 53.62 54.00 InfoBench (average) Llama2-7B-Chat 77.51 78.71 78.71 76.13 80.04 81.07 Qwen1.5-7B-Chat 80.58 79.29 79.73 78.71 81.11 81.02 Mistral-7B-Instruct 79.16 78.13 78.36 78.62 79.60 80.58 Llama2-13B-Chat 81.07 80.98 80.71 81.16 82.84 82.36 Qwen1.5-14B-Chat 82.44 80.76 80.53 79.24 83.33 83.24 Orion-14B-Chat 76.40 76.09 77.07 74.62 78.76 78.71 IFEval (loose prompt) Llama2-7B-Chat 45.84 44.18 43.99 46.95 49.17 49.17 Qwen1.5-7B-Chat 43.62 47.50 44.92 46.95 45.47 51.02 Mistral-7B-Instruct 43.25 42.51 45.84 45.66 47.69 48.24 Llama2-13B-Chat 48.06 46.77 49.17 48.24 51.39 52.31 Qwen1.5-14B-Chat 47.87 39.93 39.74 44.36 55.45 56.75 Orion-14B-Chat 37.89 36.60 40.67 38.26 53.05 50.09
@@
In DeMoRecon-Eval set, models fine-tuned on DeMoRecon-Ref outperformed those tuned on DeMoRecon-Aug on more LLMs, thanks largely to the design of our reference-based response collection process. Furthermore, larger LLMs benefit more from SFT compared to their 7B counterparts. This advantage is likely due to their greater parameter count and depth, which enable them to better capture the nuances between instruction variants and their responses, even in the absence of explicit preference learning.
However, the benefits of SFT were not as apparent across the other three benchmarks, which focus on more general instruction-following capabilities. This finding highlights a gap previously discussed in the introduction: current research often focuses on evaluating performance across challenging and diverse instructions, neglecting the subtle variations that are also crucial. DeMoRecon-Eval specifically tests instruction-following with minimal sub-instruction variance, thus revealing the unique strengths of SFT in this context.
@@
When combining DPO with SFT, significant improvements were observed across all benchmarks, including DeMoRecon-Eval. Specifically, models tuned with DeMoRecon-Ref-DPO showed enhanced performance due to the minimal contextual differences between the original and reference-generated responses. This tuning approach, which focuses on minor sub-instruction modifications, forces LLMs to prioritize the slight differences in instructions over other variables such as text quality and semantic distinctions. This focus sharpens the models’ ability to detect and adhere to nuanced instruction details.
@@
On the other three benchmarks, DPO+SFT also grants superior performance boost thanks to the preference learning. However, DeMoRecon-Ref-DPO can not grants a better performance boost on LLMs compared with DeMoRecon-Aug-DPO. This is also because of difference on the focus of evaluation. While DeMoRecon-Ref-DPO can provide better learning efficiency on following nuance difference between instructions, those three benchmarks does not exam such a capability. Thus using DeMoRecon-Ref-DPO or DeMoRecon-Aug-DPO will make less difference. On the other side, although GPT-4’s reference-based response follow instructions worse than direct response, learning from DeMoRecon-Aug-DPO does not result worse performance compared with learning from DeMoRecon-Aug-DPO.
@@
Despite these improvements, the performance boost from DeMoRecon-Ref-DPO compared to DeMoRecon-Aug-DPO was less pronounced on the other three benchmarks. This outcome suggests that the nuances DeMoRecon-Ref-DPO excels in capturing are not as critical in evaluations that do not specifically test for subtle instruction variances. Interestingly, learning from DeMoRecon-Aug-DPO did not result in poorer performance, indicating that both approaches robustly support general instruction-following capabilities even reference-based response collected from GPT-4 follows instructions worse compared with direct responding to augmented instructions.
@@
In conclusion, the experimental outcomes affirm the efficacy of DeMoRecon in enhancing LLMs’ sensitivity to subtle instructional changes, validating the value of DeMoRecon and the proposed augmentation method in advancing nuanced instruction-following capabilities. Particularly, the specially designed reference-based response collection mechanism, which adapts original responses to fit new instruction variants, has proven especially effective. Both DeMoRecon-Ref and DeMoRecon-Ref-DPO have shown to be instrumental in teaching LLMs to discern subtle differences in instructions while also enhancing their performance on common instruction-following benchmarks that focus on complex and challenging instructions.
Table 3: Results on DeMoRecon-Eval using difference fine-tuning method and number of original instructions. The base model is LLaMA-2-7B-Chat.

### v3

@@ Abstract
The effective alignment of Large Language Models (LLMs) with precise instructions is essential for their application in diverse real-world scenarios. Current methods focus on enhancing the diversity and complexity of training and evaluation samples, yet they fall short in accurately assessing LLMs’ ability to follow similar instruction variants. We introduce an effective data augmentation technique DeMoRecon that decomposes complex instructions into simpler sub-components, modifies these, and reconstructs them into new variants, thereby preserves the original instruction’s context and complexity while introducing variability, which is critical for training and evaluating LLMs’ instruction-following precision. Based on DeMoRecon, we developed the FGIV dataset which contains fine-grained instruction variants of 1,773 seed instructions to both fine-tune and evaluate LLMs. Our findings show that LLMs fine-tuned with FGIV will gain significant performance boost on both ours and commonly used instructions-following benchmarks.
@@ 1 Introduction
To address these challenges, we propose DeMoRecon, an effective data augmentation method that constructs instruction variants with only small differences in the sub-instructions of the original data by Decomposing each instruction into sub-instructions, Modifying them, and then Reconstructing these into new variants. This process preserves the original context and difficulty while increasing variability, thereby improving and assessing LLMs’ instruction-following capabilities
Based on DeMoRecon, we have developed the FGIV datasets, which Fine-Grained Instruction Variants of 1,773 seed instructions to both enhance and evaluate these capabilities. The FGIV-A dataset uses direct responses from GPT-4 based on the augmented instructions for DPO, whereas FGIV-R incorporates responses revised from the original dataset to maintain closeness to the initial instructions. Additionally, the FGIV-Eval dataset, which includes 170 instructions and their augmentations, leverages GPT-4 to assess the accuracy of following very similar instruction variants.
We leverage both instruction tuning and alignment tuning to assess the effectiveness of the FGIV datasets on LLMs and evaluating their performance across FGIV-Eval and established benchmarks like IFEval, InfoBench, and FollowBench. The results demonstrate notable enhancements in the precision with which LLMs follow instructions. Key contributions of our work include:
  • We introduce DeMoRecon, a data augmentation method that breaks down complex instructions into simpler sub-components, modifies them, and reconstructs them into new variants. This preserves the original instruction’s difficulty and context while increasing variability, thereby improving LLMs’ ability to detect nuanced differences in commands.
  • Based on DeMoRecon, we develop the FGIV dataset, which includes three components (FGIV-A, FGIV-R, and FGIV-Eval), targeting both enhanced training and rigorous evaluation of LLMs’ instruction-following abilities. This dataset is specifically designed to test LLMs on subtle variations within instructions, providing a refined tool for assessing nuanced comprehension and execution.
  • Through comprehensive testing, we show that LLMs trained with FGIV outperform those trained with traditional benchmarks in nuanced instruction-following tasks. Our findings, validated using several leading LLMs including GPT-4, establish FGIV as a critical benchmark for future assessments in this field. We have made the training data and source code fully open-source to support ongoing research.
@@ 2 Approach
Figure 1: An illustration of the proposed method of constructing FGIV.
In this section, we describe the process of constructing FGIV using DeMoRecon and detail its advantages. Figure 1 provides an illustration of our proposed instruction augmentation method, while Table 1 presents the dataset statistics for FGIV. We utilize gpt-4-0125-previewhttps://platform.openai.com/docs/models for all tasks involving large language models (LLMs). The prompt templates designed and used in our experiments are included in the Appendix.
@@ 2.1 Seed Preparation
To adequately encompass instructions of varying difficulty levels, we elected to source initial seeds from instructions developed by Xu et al. (2023) using Evol-Instructhttps://huggingface.co/datasets/WizardLMTeam/WizardLM_evol_instruct_V2_196k. We began by randomly sampling 2,000 instructions from the official dataset. Subsequently, GPT-4 was employed to exclude any instructions lacking sufficient background information. For instance, some instructions requested a summary of a passage that was not provided. After this filtering process, we established a final valid seed set, $\mathbf{S}=\{P_{i}\}_{1\leq i\leq N}$, where $N=1,773$ represents the number of valid instructions encompassing both easy and challenging topics. These seeds were then utilized to construct our FGIV.
@@ 2.2 Instruction Augmentation
Reconstruction. After modifying the sub-instruction, we reconstruct $P_{i}$ into $\bar{\mathbf{P}}_{i}^{A}=\{P_{i}\}\cup\{\bar{P}_{i}^{j}\}_{1\leq j\leq M_{i}}$, where $\bar{P}_{i}^{j}$ is formed by prompting GPT-4 with the original facts, $\mathbf{F}_{i}$, and the newly modified sub-instruction set, $\bar{\mathbf{I}}_{i}^{j}$. Specifically, in this step, we provide GPT-4 with the original instruction $P_{i}$, its sub-instructions $\mathbf{I}_{i}$, its background facts $\mathbf{F}_{i}$, and the sub-instructions $\bar{\mathbf{I}}_{i}^{j}$ where one of them are modified. We instruct GPT-4 to revise only the necessary parts of the original instruction $P_{i}$ to accommodate the changes in sub-instructions. This approach is designed to keep most of the original prompt unchanged, making the modified sub-instruction the sole variant between the original and the new instructions.
@@ 2.3 Response Collection
For evaluation, we selected a subset of 170 instructions from $S$ that each contain four sub-instructions, allowing us to control evaluation costs. We then used their augmented variants to form the test set.
For training, we construct preference pairs for preparation. For each instruction pair $(P_{i},P_{j})_{P_{i},P_{j}\in\bar{\mathbf{P}}_{i}^{A}}$, we label the response of $P_{i}$ as “good” and the response of $P_{j}$ as “bad” to create a DPO training sample. To reduce the size of the training data, we keep at most 5 pairs for each seed instructions by random sampling. We perform such process on both “Augment” dataset and “Augment with Reference” dataset to construct FGIV-A and FGIV-R.
Consequently, we construct three versions of FGIV: 1) FGIV-Eval, comprising 680 instructions designated for assessing the instruction-following capabilities of LLMs on fine-grained instruction variants; 2) FGIV-A, which include 6,445 instructions with good and bad responses collected via direct prompting of GPT-4; 3) FGIV-R, which utilizes the original instructions and their responses as references to gather responses for the instruction variants in FGIV-A.
@@ 2.4 Statistic and Analysis
FGIV- Eval Augmentation Reference FGIV-A FGIV-R Number of Instructions 680 8,705 8,705 6,445 6,445 Instruction Average Token 74 126 126 125 125 Response Average Token 466 656 631 - - $U$ of Response - 0.036 0.064 - -
Table 1: The Data statistic of FGIV.
Figure 2: The tSNE plots illustrate the semantic embeddings generated by FGIV.
Figure 2 displays the t-SNE 2D projection of the semantic embeddings of the instructions and responses in “Augment” dataset and “Augment with Reference”. In the top plot, gray points represent the original instructions in $\mathbf{S}$, and orange points represent their augmented instructions. This similarity is advantageous for LLMs during fine-tuning, as it enables them to learn to generate responses to instructions with only minor difference on sub-instructions, thereby enhancing their instruction-following capabilities.
In the lower plot, blue points denote the responses to the augmented instructions using references, while orange points signify the responses without references. The substantial overlap of blue points with gray points, which represent the responses to the original instructions, demonstrates that the responses using references closely resemble the original responses, aligning with our design objectives. Conversely, the more dispersed orange points indicate that responses without references tend to diverge from the original responses. This pattern underscores that even minimal changes to sub-instructions require LLMs to generate responses that closely match those to the original instructions. Such a task compels the LLMs to master the subtleties of handling minor differences in instruction variants, thereby demanding a high level of instruction-following capability.
Following the methodology in Guo et al. (2023), we calculate the variance of the nearest neighbor distances for the responses of “Augment” dataset and “Augment with Reference” dataset to quantify their dispersion:
@@
Here, $U$ represents the distribution uniformity, $Q$ denotes the number of responses, and $\mu$ indicates the average Euclidean distance between the semantic embedding $e_{k}$ of the $k$-th response and the context embedding $e^{\text{NN}}_{k}$ of its nearest neighbor. A larger value of $U$ implies that the responses are less uniform and thus more similar to each other. Given our goal to make the responses as similar to the original responses as possible, we expect to observe higher values of $U$ on the “Augment” dataset.
The calculated results, presented in Table 1, indicate that the distribution uniformity of the responses in “Augment with Reference” is 0.064, compared to 0.035 in “Augment”. This implies that the responses collected with reference are less uniform than those collected directly with instructions. This discrepancy further supports the observation that responses obtained using references closely resemble the original responses, more so than those obtained directly from prompting GPT-4 with only instructions.
@@ 3 Experiments
In this section, we detail the experiments conducted with LLMs fine-tuned using FGIV. We assess the accuracy of instruction-following based on evaluations from GPT-4 and rule-based (IFEval and partial FollowBench) assessments under greedy generation settings. The performance is reported across FGIV-Eval and three other well-known instruction-following benchmarks.
@@ 3.1 Baselines
With the deprecation of gpt-4-0314, a model previously central to benchmarking instruction-following across multiple datasets§§§https://platform.openai.com/docs/deprecations, previous evaluations on instruction-following benchmarks are no longer comparable, as the evaluator LLMs are no longer accessible. Our study addresses this gap by comparing baselines, official chat LLMs optimized for peak performance in various domains—including instruction-following—with models that have been further fine-tuned using the FGIV set. These official chat models utilize techniques such as instruction- and alignment-tuning, among others, to maximize performance. Unlike other studies, such as those described in (Sun et al., 2024), which primarily enhance instruction-following starting from pre-trained LLMs—a costly and time-intensive process—our research incorporates the latest updates in model versions. For all evaluations involving GPT-4, we utilize gpt-4-0125-preview.
@@
These models, available in their officially released instruction-tuned versions, are employed to precisely assess their instruction-following capabilities across various instruction variants and to explore the potential for further enhancements in this area. The specific releases utilized are detailed in Table 2 and can be accessed on Hugging Facehttps://huggingface.co.
We do not incorporate traditional text augmentation methods such as Easy Data Augmentation (EDA) (Wei and Zou, 2019) or back translation (Sennrich et al., 2016) in our study. Unlike these approaches, which entail broad textual alterations, our method specifically focuses on minor modifications to sub-instructions. This targeted approach enhances the precision of instruction-following without altering the core instructional content. Consequently, our method was not directly compared to these traditional techniques, which are also typically not considered in prior instruction-following research.
@@ 3.2 Benchmarks
Besides FGIV-Eval, we extend our performance comparison to include three of the most popular instruction-following benchmarks to further demonstrating the efficacy of our methods.
@@ 3.3 Experimental Settings
We employed both Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) on base chat models. For comparison, we also fine-tuned these models using only the seed instructions and their respective responses, with these results labeled as “seed” in Table 2.
We integrated SFT and DPO losses to fine-tune the LLMs, addressing both preference alignment and response generation, which are crucial for effective instruction-following. DPO focuses on enhancing the model’s preference for better adherence to instructions, while SFT ensures the generation of correct, precise, and high-quality responses. This combined approach has proven effective in our ablation studies. Specifically, for each preference sample, we use the “good” response to supervise the fine-tuning of LLMs alongside DPO-based learning.
@@ 3.4 Experimental Results
Model FGIV-Eval Chat Seed FGIV-A FGIV-R GPT-4 95.00 - - - GPT-4-Ref 91.03 - - - Llama2-7B-Chat 54.71 56.32 60.29 60.44 Qwen1.5-7B-Chat 61.91 71.03 72.06 70.59 Mistral-7B-Instruct 64.41 66.32 67.79 71.18 Llama2-13B-Chat 58.82 62.56 64.41 65.15 Qwen1.5-14B-Chat 73.53 72.06 79.26 81.62 Orion-14B-Chat 56.62 62.21 68.82 68.38 FollowBench (average HSR) Llama2-7B-Chat 43.10 44.49 45.56 46.41 Qwen1.5-7B-Chat 53.00 53.36 51.36 54.86 Mistral-7B-Instruct 42.55 46.56 49.63 47.68 Llama2-13B-Chat 44.94 46.56 47.23 48.07 Qwen1.5-14B-Chat 55.32 50.09 58.02 56.25 Orion-14B-Chat 44.12 48.81 53.62 54.00 InfoBench (average) Llama2-7B-Chat 77.51 78.71 80.04 81.07 Qwen1.5-7B-Chat 80.58 79.29 81.11 81.02 Mistral-7B-Instruct 79.16 78.13 79.60 80.58 Llama2-13B-Chat 81.07 80.98 82.84 82.36 Qwen1.5-14B-Chat 82.44 80.76 83.33 83.24 Orion-14B-Chat 76.40 76.09 78.76 78.71 IFEval (loose prompt) Llama2-7B-Chat 45.84 44.18 49.17 49.17 Qwen1.5-7B-Chat 43.62 47.50 45.47 51.02 Mistral-7B-Instruct 43.25 42.51 47.69 48.24 Llama2-13B-Chat 48.06 46.77 51.39 52.31 Qwen1.5-14B-Chat 47.87 39.93 55.45 56.75 Orion-14B-Chat 37.89 36.60 53.05 50.09
Table 2: Experimental results on FGIV-Eval and other instruction-following benchmarks. The performance of the official chat LLMs are recorded under “Ori”.
@@
In FGIV-Eval set, models fine-tuned on FGIV-R outperformed those tuned on FGIV-A on more LLMs, thanks largely to the design of our reference-based response collection process, which forces LLMs to prioritize the slight differences in instructions over other variables such as text quality and semantic distinctions during the fine-tuning, especially the alignment tuning part. This focus sharpens the models’ ability to detect and adhere to nuanced instruction details.
Furthermore, larger LLMs benefit more from fine-tuning compared to their 7B counterparts. This advantage is likely due to their greater parameter count and depth, which enable them to better capture the nuances between instruction variants and their responses, together with explicit preference learning.
@@
On the other three benchmarks, our method also grants superior performance boost. However, FGIV-R can not grants a better performance boost on LLMs compared with FGIV-A. This is also because of difference on the focus of evaluation. While FGIV-R can provide better learning efficiency on following nuance difference between instructions, those three benchmarks does not exam such a capability. Thus using FGIV-R or FGIV-A will make less difference. On the other side, although GPT-4’s reference-based response follow instructions worse than direct response, learning from FGIV-A does not result worse performance compared with learning from FGIV-A.
@@
Despite these improvements, the performance boost from FGIV-R compared to FGIV-A was less pronounced on the other three benchmarks. This outcome suggests that the nuances FGIV-R excels in capturing are not as critical in evaluations that do not specifically test for subtle instruction variances. Interestingly, learning from FGIV-A did not result in poorer performance, indicating that both approaches robustly support general instruction-following capabilities even reference-based response collected from GPT-4 follows instructions worse compared with direct responding to augmented instructions.
@@
In conclusion, the experimental outcomes affirm the efficacy of DeMoRecon in enhancing LLMs’ sensitivity to subtle instructional changes, validating the value of FGIV and DeMoRecon in advancing nuanced instruction-following capabilities. Particularly, the specially designed reference-based response collection mechanism, which adapts original responses to fit new instruction variants, has proven especially effective. DeMoRecon has shown to be instrumental in teaching LLMs to discern subtle differences in instructions while also enhancing their performance on common instruction-following benchmarks that focus on complex and challenging instructions, which is different from FGIV-Eval.
@@
Table 3: Results on FGIV-Eval using difference fine-tuning method and number of original instructions. The base model is LLaMA-2-7B-Chat.